A share yields the following future dividends: 40 €, 20 €, 50 €. The opportunity cost of investment, as quantified by a finance practitioner is 7%. The share can be bought today for 200 € and is going to be sold for 250 €, after 3 years. The percentage value dividend yield and the net present value of this financial investment, are:

a. 159% and 250 €

b. 73% and 180 €

c. 34% and 120 €

d. 55% and 100 €

e. 25% and 120 €


1
Expert's answer
2022-02-16T08:28:20-0500

a) Percentage dividend yield= "\frac{Annual Dividends per share}{Current Share Price}"

= "\frac{50+40+20}{200}= \frac{110}{200}" = 055%


b) "PV= \frac{D+E}{(1+r)^n}"

Where; D is any dividends expected to be paid during the period,

E is the expected stock price,

n is the number of years,

r is the real rate of return

"PV= \frac{110+250}{(1+0.07)^3}= 294"

"NPV= PV- I_0"

"294- 200= 94"